[figure: 9.7 cm]

GIANT MAGNETORESISTANCE IN THE REGION OF THE

MAGNETIC PHASE TRANSITION IN Gd_{0.5}La_{0.5}MnSi COMPOUND

S.A. Nikitin, I.A. Tskhadadze, L.A. Skipetrova, V.A. Andronova Department of Physics, Moscow State University, 119899 Moscow, Russian Federation

Gd_{0.5}La_{0.5}MnSi compound has been studied by magnetization, resistivity, and magnetoresistance measurements. As previously shown^{1} in Gd_{x}La_{1−x}MnSi series the substitution of Gd by La leads to a change of the magnetic structure from ferromagnetic (F M) type for GdMnSi to antiferromagnetic (AF M) type for LaMnSi. The temperature induced metamagnetic phase transition AF M → F M takes place for Gd_{0.5}La_{0.5}MnSi at T ≃ 100K. In the magnetically ordered region the magnetoresistance is negative, and the temperature dependence of magnetoresistance ΔR/R(T) is characterized by a sharp minimum at T ≃ 103K. The maximaum value of ΔR/R effect exceeds 3% at external magnetic field 14kOe. The absence of the saturation on ΔR/R(H) curves results from the essentially higher values of the maximaum ΔR/R effect in these compounds. The obtained results can be explained by the strong change of magnetic contribution to the conductivity due to the destruction of the layered antiferromagnetic structure. This work has been supported by the Federal Program on Support of Leading Scientific Schools N96-15-96429 and RFBR Grant N99-02-17821. 1 S.A. Nikitin, O.B. Nekrasova, T.I. Ivanova, Yu.F .Popov, R.S. Torchinova. Phys. Solid State (in Russian) 33(6) (1991) 1640.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : I.A.Tskhadadze

Address for correspondence : Department of Physics, Moscow State University, 119899 Moscow, Vorobievy Gory Russia

Email address : ioulia@rem.phys.msu.su